UNRAVELING THE INTENSITY PATTERNS IN THE INFRARED SPECTRA OF X^{−}(^{1,2}HO^{1,2}H) [X=F,Cl]

\underline{SAMANTHA HORVATH} and ANNE B. McCOY, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

The distortions from equilibrium and changes in the infrared spectra of a single water molecule when it is complexed with a halide ion (F^{−}, Cl^{−}, and Br^{−}) have been of great interest for the past several years.^{a}^{,}^{b} In these systems the bonded hydrogen stretch (ν_{2}), the so-called “proton-transfer” mode, as well as the in-plane (ν_{4}) and the out-of-plane (ν_{6}) bending modes have displayed unexpected intensity patterns in the experimental infrared spectra.^{c} An example of such anomalies can be seen in the ν_{2} mode of F^{−}(H_{2}O). This vibrational mode exhibits significant intensity in the fundamental transition, over ten times that of any other fundamental, and its frequency is shifted by nearly 1500 cm^{−1} to the red of the free OH stretch. Analysis of the first and second overtones in ν_{2} indicates that anharmonicity is not as significant as one might expect from such a large redshift of the fundamental.

We investigate the underlying motions associated with the ν_{2}, ν_{4}, and ν_{6} modes by calculating the potential energy and dipole moment surfaces at the MP2 level of theory using aug-cc-pVTZ basis sets. The X^{−}(H_{2}O) vibrational energies and wave functions are determined using one- and two-dimensional^{d} variational calculations. Our theoretical results reproduce the experimental findings extremely well and provide insight into the physics behind these intriguing spectral features.

aD. D. Kemp and M. S. Gordon, J. Phys. Chem. A \underline{109}, 7688 (2005).

bE. G. Diken et al., J. Phys. Chem. A \underline{109}, 571 (2005).

cJ. R. Roscioli, E. G. Diken, M. A. Johnson, S. Horvath, and A. B. McCoy, J. Phys. Chem. A \underline{110}, 4943 (2006).

dD. T. Colbert and W. H. Miller, J. Chem. Phys. \underline{96}, 1982 (1992).